I have attempted to solve this but have failed. Could someone Please, Please, please help me? I am new to Algebra so could you explain step by step as to how you arrived at the answer?
A ball is thrown straight up from a rooftop at 720 feet high. The formula h=-16t^2+44t+720 describes the balls height above the ground, h, in feet, t seconds after it was thrown. The ball misses the rooftop on its way down. Determine when the balls height will be 540 feet .

Answer by stanbon(75887)

A ball is thrown straight up from a rooftop at 720 feet high. The formula h=-16t^2+44t+720 describes the balls height above the ground, h, in feet, t seconds after it was thrown. The ball misses the rooftop on its way down. Determine when the balls height will be 540 feet
--------
Let the height (h) be 540. Then solve the equation for time (t).
--------
-16t^2 + 44t + 720 = 540
--------
-16t^2 + 44t + 180 = 0
-------
16t^2 - 44t - 180 = 0
-------
Use the quadratic formula to solve for time (t).
--------
t = [44 +- sqrt(44^2 - 4*16*-180)]/32
----
t = [44 +- 116]/32
----
Positive solution:
t = 5 seconds
---------------------
Cheers,
Stan H.
---------------------